Zappos Management plans to fire any Customer Loyalty Team member with an overall customer satisfaction survey rating in the 2.5%ile (bottom 2.5%). If the population of Customer Loyalty Team members have normally distributed overall customer satisfaction survey ratings with a mean µ = 4.25 and a standard deviation of σ = 0.6, what overall customer satisfaction survey ratings would result in a team member being fired? How many team members (out of N = 250) would expect to be fired?

Group of answer choices

a. Any rating of 3.05 or below; expected for 63 Customer Loyalty Team members

b. Any rating of 3.05 or below; expected for 6 Customer Loyalty Team members

c. Any rating of 3.65 or below; expected for 6 Customer Loyalty Team members

d. Any rating of 3.65 or below; expected for 63 Customer Loyalty Team members

e. Any rating of 5.45 or below; expected for 63 Customer Loyalty Team members

f. Any rating of 5.45 or below; expected for 6 Customer Loyalty Team members
Answer #1

z-value for bottom 2.5% is 1.96

x = mean + z*sigma
x = 4.25 + 1.96*0.6 = 5.426 i.e. 5.4

expected number = 250*0.025 = 6.25


f. Any rating of 5.45 or below; expected for 6 Customer Loyalty Team members